# Assessing Adversarial Robustness of Large Language Models: An Empirical Study

### v1

@@ Abstract
1 Introduction
In recent years, the field of artificial intelligence has witnessed a remarkable surge in the development and application of Large Language Models (LLMs). These models, such as ChatGPT (OpenAI, 2023a), GPT-4 (OpenAI, 2023b), and Llama-2 (Touvron et al., 2023), have demonstrated exceptional performance in various natural language understanding and generation tasks (Zhao et al., 2023a). The success of LLMs can be attributed to the innovative training techniques employed, including instruction tuning, prompt tuning, Low-Rank Adaptor (LoRA) (Hu et al., 2021; Dettmers et al., 2022). These advances have made it possible to fine-tune and infer models like Llama-2-7B on consumer-level devices, thereby increasing their accessibility and potential for integration into daily life.
@@ 1 Introduction
2 Related Work
2.1 The Evaluation of LLMs
@@
Due to the impressive performance of LLMs, there has been widespread attention to their safety and stability.  (Wang et al., 2023b) conducted an early investigation into ChatGPT and other LLMs and utilized existing benchmarks like AdvGLUE (Wang et al., 2022), ANLI (Nie et al., 2020), and DDXPlus (Tchango et al., 2022) for their evaluations. (Zhao et al., 2023b) assessed the performance of LLMs in visual inputs and their transferability to other visual-language models. On the topic of adversarial robustness,  (Wang et al., 2023a) introduced the AdvGLUE++ benchmark and proposed an approach to examine machine ethics through system prompts. (Zhu et al., 2023) proposed a unified benchmark named PromptBench to evaluate the resilience of LLMs against prompts.
@@
2.2 Robustness in NLP
With the rapid advancement in NLP research, its applications have become increasingly prevalent. This ubiquity underscores the growing need for reliable NLP systems that can effectively counteract malign content and misinformation. A seminal work (Jia & Liang, 2017) highlighted the vulnerabilities of NLP systems to adversarial attacks.
There are some works about various input perturbations, which could be categorized into three groups: character level, word level, and sentence level (Wei & Zou, 2019; Karimi et al., 2021; Ma, 2019). At the character level, adversarial attacks focus on altering individual characters within a given text.  (Ebrahimi et al., 2018) introduced HotFlip, utilizing gradient information to manipulate characters within text. (Li et al., 2019) took a different approach by identifying words and modifying their characters. At the word level, adversarial strategies revolve around replacing specific words within the content. For instance, (Alzantot et al., 2018) employs evolutionary algorithms to swap out words with their synonyms.  (Zhang et al., 2019) utilized probabilistic sampling to generate adversarial examples. Furthermore, some researchers have explored adversarial tactics at the sentence level. (Jia & Liang, 2017) suggested a method that introduces an extraneous sentence to the primary content, aiming to mislead reading models. On the other hand, (Peters et al., 2018) adopted a new approach, where they employed an encoder-decoder framework to rephrase entire sentences. Recently,  (Wang et al., 2023a) evaluated the robustness and trustworthiness of GPT-3.5 and GPT-4 models, revealing vulnerabilities such as the ease of generating toxic and biased outputs and leaking private information. Despite GPT-4’s improved performance on standard benchmarks, it is more susceptible to adversarial prompts, highlighting the need for rigorous trustworthiness guarantees and robust safeguards against new adaptive attacks.
@@ 2.2 Robustness in NLP
3 Preliminaries
3.1 Open-source Large Language Models
@@
Table 1: The open-source large language models in our study
3.2 Fine-tuning Techniques
@@
4 Methods
Figure 1: The framework of our adversarial robustness assessment
4.1 Adversarial Attack
@@
4.2 Geometry Attack Methodology
In our research, we extend the basic principles of adversarial attacks in the context of LLMs. Our focus is on exploiting geometric attacks (Meng & Wattenhofer, 2020; Meng et al., 2022) to assess the vulnerability of LLMs to adversarial perturbations. We propose a systematic methodology grounded in geometric attack insights. The following sections detail the steps of our approach:
@@ 4.2 Geometry Attack Methodology
5 Experiment Settings
5.1 Experiment Pipeline
@@
5.2 Datasets
To evaluate the model’s performance under various tasks and its resilience to attacks, we employed five classification datasets, categorized into binary and multi-class classifications. For binary classification, the datasets include IMDB (Maas et al., 2011), MRPC (Dolan & Brockett, 2005), and SST-2 (Socher et al., 2013), and for multiclass classification, AGNews (Zhang et al., 2015) and DBpedia (Auer et al., 2007) are used. We will provide a more detailed introduction to these tasks/datasets in Appendix A.2
5.3 Evaluation Metrics
@@
Table 2: The evaluation metrics
6 Experimental Results
@@ 6.1 Model Size (RQ1)
IMDB SST-2 MRPC
Acc Acc/attack $\uparrow$ ASR $\downarrow$ Replacement Acc Acc/attack $\uparrow$ ASR $\downarrow$ Replacement Acc Acc/attack $\uparrow$ ASR $\downarrow$ Replacement
T5-60m 0.8484 0.1256 0.8491 0.0929 0.9083 0.2419 0.7304 0.1428 0.8048 0.0325 0.9594 0.0720
T5-220m 0.8011 0.0436 0.9463 0.0722 0.8884 0.1228 0.8622 0.1611 0.8035 0.0633 0.9203 0.1100
T5-770m 0.9048 0.1536 0.8312 0.1143 0.8739 0.0534 0.9395 0.1785 0.8924 0.1992 0.7771 0.1203
T5-3b 0.9146 0.3259 0.6436 0.1413 0.9495 0.1563 0.8437 0.1987 0.8584 0.1074 0.8739 0.0917
T5-11b 0.9122 0.3098 0.6604 0.1330 0.9656 0.2248 0.7672 0.2043 0.8877 0.0712 0.9196 0.0873
OPT-125m 0.8616 0.6637 0.2297 0.0365 0.8807 0.7409 0.1587 0.0607 0.8321 0.6504 0.2184 0.0332
OPT-350m 0.8564 0.6924 0.1915 0.0305 0.9011 0.7716 0.1437 0.0598 0.8956 0.6741 0.2473 0.0437
OPT-1.3b 0.9231 0.7515 0.1859 0.0421 0.9443 0.8227 0.1288 0.0733 0.9134 0.7721 0.1547 0.0419
OPT–2.7b 0.9198 0.7651 0.1682 0.0396 0.8897 0.7921 0.1097 0.0476 0.9128 0.7854 0.1396 0.0533
OPT-6.7b 0.9408 0.7864 0.1641 0.0528 0.9693 0.8682 0.1043 0.0550 0.9096 0.7902 0.1313 0.0579
OPT-13b 0.9431 0.8016 0.1500 0.0671 0.9656 0.7867 0.1853 0.0719 0.9254 0.8183 0.1157 0.0560
Llama-7b 0.9483 0.8203 0.1350 0.0816 0.9683 0.8203 0.1528 0.0916 0.9277 0.8256 0.1101 0.0637
Llama-13b 0.9472 0.8237 0.1304 0.0875 0.9632 0.8124 0.1566 0.0877 0.9198 0.8107 0.1186 0.0742
Table 3: IMDB, SST-2 and MRPC Results
@@
AGNews DBPedia
Acc Acc/attack $\uparrow$ ASR $\downarrow$ Replacement Acc Acc/attack $\uparrow$ ASR $\downarrow$ Replacement
T5-60m 0.8606 0.3608 0.5807 0.1740 0.9817 0.3974 0.5952 0.1187
T5-220m 0.9084 0.5370 0.4098 0.1864 0.9765 0.4082 0.5819 0.1234
T5-770m 0.9278 0.6597 0.2896 0.1860 0.9921 0.7476 0.2464 0.1483
T5-3b 0.9193 0.7267 0.2102 0.1834 0.9914 0.8608 0.1317 0.1550
T5-11b 0.9212 0.8469 0.1531 0.1867 0.9919 0.8815 0.1113 0.1724
OPT-125m 0.8152 0.6040 0.2591 0.0809 0.9034 0.6512 0.2792 0.0534
OPT-350m 0.8321 0.6036 0.2746 0.0864 0.9511 0.6718 0.2937 0.0305
OPT-1.3b 0.8806 0.6316 0.2828 0.0912 0.9784 0.7046 0.2798 0.0496
OPT–2.7b 0.9175 0.7028 0.2340 0.0833 0.9822 0.7513 0.2351 0.0552
OPT-6.7b 0.9341 0.7143 0.2353 0.0756 0.9907 0.7780 0.2147 0.0641
OPT-13b 0.9456 0.7745 0.1809 0.0941 0.9916 0.7912 0.2021 0.0576
Llama-7b 0.9328 0.7315 0.2158 0.0864 0.9908 0.7596 0.2333 0.0881
Llama-13b 0.9338 0.7688 0.1767 0.0837 0.9921 0.7286 0.2656 0.0912
Table 4: AGNews and DBPedia Results
In Tables 4, analyzing results from multi-class classification tasks, a distinct pattern emerges. These datasets reveal enhanced stability against synonym substitution attacks. For T5 models, the data shows a lower ASR on these tasks compared to binary datasets, suggesting a better resistance to attacks in complex classification scenarios. In contrast, OPT and Llama models exhibit a higher ASR on the AGNews and DBpedia14 datasets. Another result is that for both T5 and OPT series, there is a marked decline in ASR around the 770 million or 1 billion parameter threshold. This indicates an increased robustness and better handling of adversarial attacks with the scale-up of model size.
6.1.1 Analysis
@@
Figure 2: The experimental results of different models on various datasets.
6.2 LLMs Fine-tuning Techniques (RQ2)
6.2.1 Instruction Tuning
@@
Table 5: Experimental results comparing T5 and Flan-T5 after instruction tuning using the IMDB dataset.
Figure 3: The experimental results of T5 and Flan-T5 on IMDB dataset
(a) Different precisions on the IMDB dataset (T5 model)
(b) Different precisions on the IMDB dataset (OPT Model)
(c) Results of LoRA on IMDB dataset
Figure 4: Precision and attack success rate
6.2.2 Precisions
@@
Table 6: Results of T5-770m and OPT-1.3b models under different precision settings, including fp16, int8 and int4. The performance is evaluated with IMDB dataset.
@@
Table 7: Results of the T5-770m, OPT-1.3b, and OPT-2.7b models’ performance with and without the application of LoRA, using the IMDB dataset.
6.2.3 LoRA
@@ 6.3 Model Architectures (RQ3)
Table 8: Experimental results of OPT models with/without classification head with IMDB dataset.
Figure 5: Results of classfication head on IMDB dataset
7 Conclusion
Ethics statement
Appendix A Appendix
A.1 Experiments Setup
A.2 Dataset
The statistics of these five datasets are presented in Table 9.
Table 9: Statistics of the Datasets
Dataset Labels Avg Length Train Test
IMDB 2 279.48 25000 25000
MRPC 2 52.89 3670 1730
SST-2 2 19.81 67300 1820
AGNews 4 43.93 120000 7600
DBPedia 14 55.14 560000 7000
A.3 Attack Examples
In this section, we are going to show some adversarial examples for different tasks and models.
Table 10: Adversarial Examples for Various Tasks and Models

### v2

@@ Abstract.
Evaluation, Robustness, Adversarial Attack, Scaling Law, LLMs Fine-tuning
ccs: Computing methodologies Natural language processing
1. Introduction
In recent years, the field of artificial intelligence has witnessed a remarkable surge in the development and application of Large Language Models (LLMs). These models, such as ChatGPT (OpenAI, 2023b), GPT-4 (OpenAI, 2023a), and Llama-2 (Touvron et al., 2023), have demonstrated exceptional performance in various natural language understanding and generation tasks (Zhao et al., 2023b). The success of LLMs can be attributed to the innovative training techniques employed, including instruction tuning, prompt tuning, Low-Rank Adaptor (LoRA) (Hu et al., 2021; Dettmers et al., 2022). These advances have made it possible to fine-tune and infer models like Llama-2-7B on consumer-level devices, thereby increasing their accessibility and potential for integration into daily life.
@@ 1. Introduction
2. Related Work
2.1. The Evaluation of LLMs
@@
Due to the impressive performance of LLMs, there has been widespread attention to their safety and stability.  (Wang et al., 2023b) conducted an early investigation into ChatGPT and other LLMs and utilized existing benchmarks like AdvGLUE (Wang et al., 2022), ANLI (Nie et al., 2020), and DDXPlus (Tchango et al., 2022) for their evaluations. (Zhao et al., 2023a) assessed the performance of LLMs in visual inputs and their transferability to other visual-language models. On the topic of adversarial robustness,  (Wang et al., 2023a) introduced the AdvGLUE++ benchmark and proposed an approach to examine machine ethics through system prompts. (Zhu et al., 2023) proposed a unified benchmark named PromptBench to evaluate the resilience of LLMs against prompts.
@@
2.2. Robustness in NLP
With the rapid advancement in NLP research, its applications have become increasingly prevalent. This ubiquity underscores the growing need for reliable NLP systems that can effectively counteract malign content and misinformation. A seminal work (Jia and Liang, 2017) highlighted the vulnerabilities of NLP systems to adversarial attacks.
There are some works about various input perturbations, which could be categorized into three groups: character level, word level, and sentence level (Wei and Zou, 2019; Karimi et al., 2021; Ma, 2019). At the character level, adversarial attacks focus on altering individual characters within a given text.  (Ebrahimi et al., 2018) introduced HotFlip, utilizing gradient information to manipulate characters within text. (Li et al., 2019) took a different approach by identifying words and modifying their characters. At the word level, adversarial strategies revolve around replacing specific words within the content. For instance, (Alzantot et al., 2018) employs evolutionary algorithms to swap out words with their synonyms.  (Zhang et al., 2019) utilized probabilistic sampling to generate adversarial examples. Furthermore, some researchers have explored adversarial tactics at the sentence level. (Jia and Liang, 2017) suggested a method that introduces an extraneous sentence to the primary content, aiming to mislead reading models. On the other hand, (Peters et al., 2018) adopted a new approach, where they employed an encoder-decoder framework to rephrase entire sentences. Recently,  (Wang et al., 2023a) evaluated the robustness and trustworthiness of GPT-3.5 and GPT-4 models, revealing vulnerabilities such as the ease of generating toxic and biased outputs and leaking private information. Despite GPT-4’s improved performance on standard benchmarks, it is more susceptible to adversarial prompts, highlighting the need for rigorous trustworthiness guarantees and robust safeguards against new adaptive attacks.
@@ 2.2. Robustness in NLP
3. Preliminaries
3.1. Open-source Large Language Models
@@
Table 1. The open-source large language models in our study
3.2. Fine-tuning Techniques
@@
4. Methods
Figure 1. The framework of our adversarial robustness assessment
4.1. Adversarial Attack
@@
4.2. Geometry Attack Methodology
In our research, we extend the basic principles of adversarial attacks in the context of LLMs. Our focus is on exploiting geometric attacks (Meng and Wattenhofer, 2020; Meng et al., 2022) to assess the vulnerability of LLMs to adversarial perturbations. We propose a systematic methodology grounded in geometric attack insights. The following sections detail the steps of our approach:
@@ 4.2. Geometry Attack Methodology
5. Experiment Settings
5.1. Experiment Pipeline
@@
5.2. Datasets
To evaluate the model’s performance under various tasks and its resilience to attacks, we employed five classification datasets, categorized into binary and multi-class classifications. For binary classification, the datasets include IMDB (Maas et al., 2011), MRPC (Dolan and Brockett, 2005), and SST-2 (Socher et al., 2013), and for multiclass classification, AGNews (Zhang et al., 2015) and DBpedia (Auer et al., 2007) are used. We will provide a more detailed introduction to these tasks/datasets in Appendix A.2
5.3. Evaluation Metrics
@@
Table 2. The evaluation metrics
6. Experimental Results
@@ 6.1. Model Size (RQ1)
Acc Acc/attack $\uparrow$ ASR $\downarrow$ Replacement
T5-60m 0.8484 0.1256 0.8491 0.0929
T5-220m 0.8011 0.0436 0.9463 0.0722
T5-770m 0.9048 0.1536 0.8312 0.1143
T5-3b 0.9146 0.3259 0.6436 0.1413
T5-11b 0.9122 0.3098 0.6604 0.1330
OPT-125m 0.8616 0.6637 0.2297 0.0365
OPT-350m 0.8564 0.6924 0.1915 0.0305
OPT-1.3b 0.9231 0.7515 0.1859 0.0421
OPT–2.7b 0.9198 0.7651 0.1682 0.0396
OPT-6.7b 0.9408 0.7864 0.1641 0.0528
OPT-13b 0.9431 0.8016 0.1500 0.0671
Llama-7b 0.9483 0.8203 0.1350 0.0816
Llama-13b 0.9472 0.8237 0.1304 0.0875
Table 3. IMDB Dataset Results
Acc Acc/attack $\uparrow$ ASR $\downarrow$ Replacement
T5-60m 0.9083 0.2419 0.7304 0.1428
T5-220m 0.8884 0.1228 0.8622 0.1611
T5-770m 0.8739 0.0534 0.9395 0.1785
T5-3b 0.9495 0.1563 0.8437 0.1987
T5-11b 0.9656 0.2248 0.7672 0.2043
OPT-125m 0.8807 0.7409 0.1587 0.0607
OPT-350m 0.9011 0.7716 0.1437 0.0598
OPT-1.3b 0.9443 0.8227 0.1288 0.0733
OPT–2.7b 0.8897 0.7921 0.1097 0.0476
OPT-6.7b 0.9693 0.8682 0.1043 0.0550
OPT-13b 0.9656 0.7867 0.1853 0.0719
Llama-7b 0.9683 0.8203 0.1528 0.0916
Llama-13b 0.9632 0.8124 0.1566 0.0877
Table 4. SST-2 Dataset Results
Acc Acc/attack $\uparrow$ ASR $\downarrow$ Replacement
T5-60m 0.8048 0.0325 0.9594 0.0720
T5-220m 0.8035 0.0633 0.9203 0.1100
T5-770m 0.8924 0.1992 0.7771 0.1203
T5-3b 0.8584 0.1074 0.8739 0.0917
T5-11b 0.8877 0.0712 0.9196 0.0873
OPT-125m 0.8321 0.6504 0.2184 0.0332
OPT-350m 0.8956 0.6741 0.2473 0.0437
OPT-1.3b 0.9134 0.7721 0.1547 0.0419
OPT–2.7b 0.9128 0.7854 0.1396 0.0533
OPT-6.7b 0.9096 0.7902 0.1313 0.0579
OPT-13b 0.9254 0.8183 0.1157 0.0560
Llama-7b 0.9277 0.8256 0.1101 0.0637
Llama-13b 0.9198 0.8107 0.1186 0.0742
Table 5. MRPC Dataset Results
@@
Acc Acc/attack $\uparrow$ ASR $\downarrow$ Replacement
T5-60m 0.8606 0.3608 0.5807 0.1740
T5-220m 0.9084 0.5370 0.4098 0.1864
T5-770m 0.9278 0.6597 0.2896 0.1860
T5-3b 0.9193 0.7267 0.2102 0.1834
T5-11b 0.9212 0.8469 0.1531 0.1867
OPT-125m 0.8152 0.6040 0.2591 0.0809
OPT-350m 0.8321 0.6036 0.2746 0.0864
OPT-1.3b 0.8806 0.6316 0.2828 0.0912
OPT–2.7b 0.9175 0.7028 0.2340 0.0833
OPT-6.7b 0.9341 0.7143 0.2353 0.0756
OPT-13b 0.9456 0.7745 0.1809 0.0941
Llama-7b 0.9328 0.7315 0.2158 0.0864
Llama-13b 0.9338 0.7688 0.1767 0.0837
Table 6. AGNews Dataset Results
Acc Acc/attack $\uparrow$ ASR $\downarrow$ Replacement
T5-60m 0.9817 0.3974 0.5952 0.1187
T5-220m 0.9765 0.4082 0.5819 0.1234
T5-770m 0.9921 0.7476 0.2464 0.1483
T5-3b 0.9914 0.8608 0.1317 0.1550
T5-11b 0.9919 0.8815 0.1113 0.1724
OPT-125m 0.9034 0.6512 0.2792 0.0534
OPT-350m 0.9511 0.6718 0.2937 0.0305
OPT-1.3b 0.9784 0.7046 0.2798 0.0496
OPT–2.7b 0.9822 0.7513 0.2351 0.0552
OPT-6.7b 0.9907 0.7780 0.2147 0.0641
OPT-13b 0.9916 0.7912 0.2021 0.0576
Llama-7b 0.9908 0.7596 0.2333 0.0881
Llama-13b 0.9921 0.7286 0.2656 0.0912
Table 7. DBPedia Dataset Results
In Tables 6 and Tables  7, analyzing results from multi-class classification tasks, a distinct pattern emerges. These datasets reveal enhanced stability against synonym substitution attacks. For T5 models, the data shows a lower ASR on these tasks compared to binary datasets, suggesting a better resistance to attacks in complex classification scenarios. In contrast, OPT and Llama models exhibit a higher ASR on the AGNews and DBpedia14 datasets. Another result is that for both T5 and OPT series, there is a marked decline in ASR around the 770 million or 1 billion parameter threshold. This indicates an increased robustness and better handling of adversarial attacks with the scale-up of model size.
6.1.1. Analysis
@@
Figure 2. The experimental results of different models on various datasets.
6.2. LLMs Fine-tuning Techniques (RQ2)
6.2.1. Instruction Tuning
@@
Table 8. Experimental results comparing T5 and Flan-T5 after instruction tuning using the IMDB dataset.
Figure 3. The experimental results of T5 and Flan-T5 on IMDB dataset
Figure 4. Different precisions on the IMDB dataset (T5 model)
Figure 5. Different precisions on the IMDB dataset (OPT Model)
Figure 6. Results of LoRA on IMDB dataset
Figure 7. Comparison of different models and precisions on the IMDB dataset
6.2.2. Precisions
@@
Table 9. Results of T5-770m and OPT-1.3b models under different precision settings, including fp16, int8 and int4. The performance is evaluated with IMDB dataset.
@@
Table 10. Results of the T5-770m, OPT-1.3b, and OPT-2.7b models’ performance with and without the application of LoRA, using the IMDB dataset.
6.2.3. LoRA
@@ 6.3. Model Architectures (RQ3)
Table 11. Experimental results of OPT models with/without classification head with IMDB dataset.
Figure 8. Results of classfication head on IMDB dataset
6.4. Attack Examples
In this section, we are going to show some adversarial examples for different tasks and models in Table  12
Table 12. Adversarial Examples for Various Tasks and Models
@@ 6.4. Attack Examples
7. Conclusion
Ethics statement
Appendix A Research Methods
A.1. Experiments Setup
A.2. Dataset
The statistics of these five datasets are presented in Table 13.
Table 13. Statistics of the Datasets
Dataset Labels Avg Length Train Test
IMDB 2 279.48 25000 25000
MRPC 2 52.89 3670 1730
SST-2 2 19.81 67300 1820
AGNews 4 43.93 120000 7600
DBPedia 14 55.14 560000 7000